# Towards Implicit Aggregation: Robust Image Representation for Place Recognition in the Transformer Era

### v1

@@ 4.2 Implementation Details
The experiments are conducted on the NVIDIA RTX A6000 GPU using PyTorch. We use DINOv2-base-register as the backbone and only fine-tune the last four transformer blocks with the previous layers frozen. The token dimension in backbone is 768, and the number of our aggregation tokens is 8, thus outputting 6144-dim global descriptors. The image resolution is $\times224$ for training and $\times322$ for inference, as in SALAD salad. We employ the multi-similarity loss multiloss for training, with hyperparameters set following the GSV-Cities work gsv. The model is trained using the Adam optimizer with an initial learning rate of 0.00005, halved every 3 epochs. Each training batch contains 120 places, with 4 images per place (i.e., 480 images). Besides, we set the maximum epochs to 20.
@@ 4.3 Comparisons with State-of-the-Art Methods
This section shows the experimental comparison of our ImAge with SOTA methods, including 11 single-stage VPR methods: NetVLAD netvlad, SFRS sfrs, CosPlace cosplace, MixVPR mixvpr, EigenPlaces eigenplaces, CricaVPR cricavpr, SALAD salad, SALAD-CM salad-cm, BoQ boq, SuperVLAD SuperVLAD and EDTformer edtformer, as well as 2 two-stage VPR methods (TransVPR transvpr and SelaVPR selavpr) that leverage local features for re-ranking. The latest studies, CricaVPR, SALAD, SALAD-CM, BoQ, SelaVPR, SuperVLAD, and EDTformer, all use the foundation model DINOv2 as the backbone to extract deep features and achieve SOTA results. Our method mainly adopts DINOv2-base-register in experiments. Additionally, Cosplace and EigenPlaces construct an extra large-scale dataset (SF-XL) for training. CircaVPR, SALAD, BoQ, and EDTformer are trained on the GSV-Cities dataset, while SALAD-CM combines GSV-Cities and MSLS-train for training. Our work further merges Pitts30k-train, MSLS-train, SF-XL, and GSV-Cities for training, following the process in SelaVPR++ selavpr++. Table 2 presents the comprehensive quantitative results. Moreover, to enable a fairer comparison among three leading aggregation methods (NetVLAD, SALAD, and BoQ) and our ImAge, we conduct a consistent comparison using the same setup (backbone, training data, image resolution), as shown in Table 3. The experiments using other transformer backbones (ViT and CLIP) are shown in Appendix D.
@@
NetVLAD netvlad 32768 81.9 91.2 93.7 53.1 66.5 71.1 35.1 47.4 51.7 60.6 68.9 74.6 6.4 10.1 12.5
SFRS sfrs 4096 89.4 94.7 95.9 69.2 80.3 83.1 41.6 52.0 56.3 81.0 88.3 92.4 16.1 23.9 28.4
TransVPR transvpr / 89.0 94.9 96.2 86.8 91.2 92.4 63.9 74.0 77.5 79.0 82.2 85.1 63.5 68.5 70.2
CosPlace cosplace 512 88.4 94.5 95.7 82.8 89.7 92.0 61.4 72.0 76.6 81.9 90.2 92.7 58.5 73.7 79.4
MixVPR mixvpr 4096 91.5 95.5 96.3 88.0 92.7 94.6 64.0 75.9 80.6 85.1 91.7 94.3 76.2 86.9 90.3
EigenPlaces eigenplaces 2048 92.5 96.8 97.6 89.1 93.8 95.0 67.4 77.1 81.7 93.0 96.2 97.5 71.2 83.8 88.1
SelaVPR selavpr / 92.8 96.8 97.7 90.8 96.4 97.2 73.5 87.5 90.6 94.0 96.8 97.5 87.3 93.8 95.6
CricaVPR cricavpr 4096 94.9 97.3 98.2 90.0 95.4 96.4 69.0 82.1 85.7 93.0 97.5 98.1 90.7 96.3 97.6
SuperVLAD SuperVLAD 3072 95.0 97.4 98.2 92.2 96.6 97.4 75.3 86.8 89.9 95.2 97.8 98.1 91.0 96.4 97.7
SALAD salad 8448 92.5 96.4 97.5 92.2 96.4 97.0 75.0 88.8 91.3 94.6 97.5 97.8 89.7 95.5 97.0
SALAD-CM salad-cm 8448 92.7 96.8 97.9 94.2 97.2 97.4 82.7 91.2 92.7 96.8 97.5 97.8 96.0 98.5 99.2
BoQ boq 12288 93.7 97.1 97.9 93.8 96.8 97.0 79.0 90.3 92.0 96.5 97.8 98.4 90.6 96.0 97.5
EDTformer edtformer 4096 93.4 97.0 97.9 92.0 96.6 97.2 78.4 89.8 91.9 97.1 98.1 98.4 88.3 95.3 97.0
@@
Besides, we also conduct the consistent comparison experiments on some supplementary datasets, including Baidu Mall baidumall, SPED SPED, Pitts250k pitts, St. Lucia stlucia, and SVOX SVOX, and the results are shown in Table 4. Compared to the three SOTA explicit aggregation methods, our ImAge achieves the best R@1 performance on all supplementary datasets. In particular, on Baidu Mall, which is the only indoor dataset and exhibits a distinct visual distribution from the other outdoor datasets, our method achieves the best performance, outperforming NetVLAD, SALAD, and BoQ with 0.8%, 3.3%, and 5.0% absolute R@1 improvements, respectively. This demonstrates that the global descriptors produced by our ImAge method through implicit aggregation are not only highly robust against common visual changes but also exhibit superior generalization ability.
@@ 4.4 Ablation Studies
Effect of aggregation tokens initialization. To validate the effectiveness of our proposed initialization methods for agg tokens, we conduct an ablation study using four initialization strategies: zero initialization (i.e., no initialization), normal distribution initialization (commonly used for the class token or register tokens initialization dinoregisters), vanilla cluster centers (yielded by $k$-means) initialization, and L2-normalized cluster centers initialization (i.e., ours). We consistently use 8 agg tokens and prepend them to the patch tokens before the fourth-to-last transformer block. The experimental results are presented in Table 6. Zero initialization produces uniform representations at the beginning, limiting (even harming) interaction between agg tokens and patch features, and hindering global context modeling. In contrast, normal initialization provides a better inductive bias during early training and introduces slight randomness into the agg tokens, which helps break symmetry to get better performance. However, both initialization methods lack any visual prior, forcing the agg tokens to learn the patterns relevant to VPR from scratch, which constrains their final performance. Initializing agg tokens with cluster centers can be viewed as injecting a data-driven prior. These centers, obtained via unsupervised clustering of descriptors from randomly sampled training images, tend to capture common visual patterns. Such initialization can facilitate agg tokens to learn meaningful global information and diminish useless elements. Moreover, L2-normalized cluster centers offer more robust initializations for agg tokens by mitigating the influence of outliers, thereby achieving the optimal performance on all datasets.
@@ Appendix B Limitations and Future Work
While our study provides a novel insight (i.e., implicit aggregation without an explicit aggregator) into achieving robust global image representation for VPR, we acknowledge three limitations of our work: 1) Although our ImAge method demonstrates universality for the transformer-based models, compared to the foundation models pre-trained on the massive dataset (e.g., DINOv2 pre-trained on the LVD-142M dataset dinov2), the performance improvement is less pronounced when using backbones without pre-training on sufficiently massive data (e.g., the ViT pre-trained only on ImageNet ImageNet). This will be shown in Appendix D. However, this also suggests that with the advancement of increasingly powerful foundation models, the superiority of our approach compared to existing VPR methods may become more prominent. 2) The proposed method may not be a good choice when we want to keep the backbone frozen (like AnyLoc anyloc) or when the backbone is extremely expensive to fine-tune. However, it is also worth noting that we only fine-tune the last few blocks of the transformer backbone in most cases, which is relatively cheap. 3) Although this work focuses on the VPR task, we believe that the proposed ImAge is broadly applicable to a wide range of image (or other modalities) retrieval tasks. The potential of our ImAge method for more applications in the machine learning community needs to be further explored through more experiments in future work.
@@ Appendix C More Details about the Relations & Differences to Other Methods
Our ImAge design draws inspiration from prior works, particularly DINOv2-register dinoregisters and prompt tuning prompttuning. Both approaches introduce additional tokens to the transformer backbone before the first encoder block. However, their objectives and usage differ fundamentally from ours. Prompt Tuning aims to adapt a frozen model to downstream tasks by learning a set of prompt tokens in a parameter-efficient manner, while these tokens are typically excluded from the final representation. DINOv2-register introduces additional register tokens to mitigate artifacts in the feature maps. Although the study shows that these registers may capture certain global information, they are ultimately discarded, and only the patch and class tokens are used for downstream tasks. In contrast, our ImAge method introduces the aggregation (agg) tokens before a particular transformer block, and utilizes agg tokens from the output of the last block as the final global image representation, which provides a reverse perspective compared to these approaches. In addition, while the class token within the transformer backbone is sometimes used as a global representation, our ImAge differs in several key aspects and demonstrates superior scalability. First, the class token is typically introduced at the beginning of the transformer and starts to learn from shallow features, which may limit its flexibility and make it difficult to fully capture task-specific complex semantics. Second, the class token is a single fixed embedding, which inherently restricts its representational capacity. Although it may suffice for relatively simple classification tasks, it often proves inadequate for more complex scenarios requiring richer and more flexible representations. In contrast, our ImAge introduces agg tokens with customizable positions and quantities, allowing them to fully learn task-specific global features. Moreover, we also design a tokens initialization method based on the $k$-means algorithm, which is significantly different from other works. Additionally, among existing VPR methods, BoQ boq also offers some valuable insights for our work. However, it elaborately designs an explicit aggregator consisting of additional encoder blocks and cross-attention layers, which aggregates global information from patch tokens into a set of extra learnable queries. In contrast, we focus on the transformer backbone itself and make use of the inherent self-attention mechanism. Our study reveals a new insight: the aggregation function, previously implemented through an exquisitely designed aggregator, already appears naturally in the transformer backbone. We demonstrate that, by adding just some additional tokens, we can fully develop this implicit and progressive aggregation behavior.
@@ Appendix D Comparison to NetVLAD Using Other Transformer Backbones
In the main paper, we conduct experiments using the DINOv2 backbone to validate the effectiveness of our method. Notably, DINOv2 is a foundation model based on the ViT architecture and pre-trained on the large-scale curated LVD-142M dataset. However, our method is also applicable to other transformer models. To this end, we conduct additional experiments using the CLIP clip model and a ViT model pre-trained only on ImageNet. The results are shown in Table 8. We observe that our proposed method, ImAge, consistently achieves higher R@1 performance across all datasets compared to the explicit aggregation method NetVLAD. However, the performance gains of ImAge are less pronounced (except for ViT-ImAge on MSLS-val) compared to using DINOv2-base-register as the backbone. This observation aligns with the prior study eomt, which suggests that foundation models pre-trained on large-scale datasets (significantly larger than ImageNet) are more capable of utilizing additional tokens to capture global information. Moreover, using CLIP as the backbone yields significantly less improvement than DINOv2. Although CLIP is a widely used foundation model (i.e., a vision-language model), its pre-training data and objectives differ considerably from those of the VPR task, making it not a promising choice. This is consistent with the prior work AnyLoc anyloc, which suggests that CLIP performs significantly worse than DINOv2 in outdoor VPR scenarios.
Table 8: Results of NetVLAD and our ImAge using CLIP (base version, only vision encoder) and ViT (base version) as backbone. All models are trained on the GSV-Cities dataset with the batch size equal to 120. The learning rate is 0.00006 for the CLIP-based model and 0.0003 for the ViT-based model. For ViT, the last two blocks are directly truncated and all other blocks are trainable, as in benchmark; SuperVLAD. For CLIP, we only train the last 6 blocks with the previous layers frozen. All methods produce 768*8-dimensional descriptors, i.e., 8 clusters for NetVLAD and 8 aggregation tokens for ImAge, the same as in the main paper.
@@ Appendix E Improving Other VPR Methods with ImAge
Since our ImAge is a general image representation method for VPR, it can not only be implemented based on different transformer backbones, but also can be combined with some other VPR methods to improve their performance. This section uses the CricaVPR cricavpr method as an example to conduct experiments, and the results are shown in Table 9. It can be seen that our method significantly improves the performance.
@@ Appendix I More Details about Datasets
The testing datasets used in our experiments, including Pitts30k, Pitts250k, Tokyo24/7, Nordland, SPED, St. Lucia, and SVOX, are organized following the Visual Geo-localization (VG) benchmark benchmark. Notably, we use the official version MSLS dataset as in previous work msls; mixvpr; cricavpr; boq; salad. This version of MSLS-val only consists of 740 query images, which is different from the version in the VG benchmark benchmark. In addition, there are also several versions of the Nordland dataset in the VPR task. In our experiments, we use the version in the VG benchmark benchmark, which employs the summer images as the database and the winter images as queries, each containing 27592 images. Baidu Mall baidumall is a well-known indoor dataset for image-based localization. All images are collected at a shopping mall that is over 5000 square meters with many challenging elements, such as transparent windows, reflective materials, repetitive structures, dynamic pedestrians, etc.
Moreover, in the comprehensive comparison (i.e., Table 2 in main paper) with other SOTA methods, we merge Pitts30k-train, MSLS-train, SF-XL, and GSV-Cities for training, following the approach in SelaVPR++ selavpr++. Specifically, we process datasets other than GSV-Cities to divide places into a finite number of categories, thus facilitating fully supervised training with the multi-similarity loss multiloss.
@@
NetVLAD netvlad and SFRS sfrs both consist of a VGG16 backbone and a NetVLAD aggregator, and use Pitts30k as the training dataset. The latter employs self-supervised image-to-region similarities to mine hard positive samples for training a more robust model. In our experiments, we use their PyTorch implementations for comparison.
CosPlace cosplace and EigenPlaces eigenplaces both frame VPR training as a classification task and use the SF-XL dataset to train their models. For Cosplace, we use the official model based on the VGG16 backbone (with the 512-dim output feature) for testing. For EigenPlaces, we utilize its official implementation and the best configuration based on the ResNet50 backbone to output 2048-dim global descriptors for comparison.
MixVPR mixvpr aggregates the deep features using the multi-layer perceptrons and trains the model with multi-similarity loss multiloss on the GSV-Cities gsv dataset. We apply its best-performing configuration (ResNet50 with 4096-dim output features) for comparison.
CricaVPR cricavpr, SuperVLAD SuperVLAD, SALAD salad, BoQ boq, and EDTformer edtformer all use the foundation model DINOv2-base dinov2 as the backbone to extract deep features, and train their models on GSV-Cities with the multi-similarity loss. In the comparison experiments, we consistently use their official implementations and the best configurations.
SALAD-CM salad-cm is an improvement of SALAD. This work analyzes the Geographic Distance Sensitivity of VPR embeddings and proposes a novel mining strategy to address it. Moreover, SALAD-CM first trains the model using both the GSV-Cities and MSLS datasets for better performance. In the comparison experiments, we follow its official implementation.
The rest TransVPR transvpr and SelaVPR selavpr are two-stage VPR methods. These works provide two models: one trained for testing on urban datasets (e.g., Pitts30k and Tokyo24/7), and another trained for testing on datasets that may contain suburban and natural scenes (e.g., MSLS and Nordland). We follow the usage in their original paper for comparison experiments.
@@
Fig. 8 is the snapshot of the MSLS place recognition challenge msls leaderboard at the time of submission, and our ImAge method ranks 1st.

### v2

@@ 4.2 Implementation Details
The experiments are conducted on the NVIDIA RTX A6000 GPU using PyTorch. We use DINOv2-base-register as the backbone and only fine-tune the last four transformer blocks with the previous layers frozen. The token dimension in backbone is 768, and the number of our aggregation tokens is 8, thus outputting 6144-dim global descriptors. The image resolution is $\times224$ for training and $\times322$ for inference, as in SALAD Izquierdo and Civera (2024b). We employ the multi-similarity loss Wang et al. (2019) for training, with hyperparameters set following the GSV-Cities work Ali-bey et al. (2022). The model is trained using the Adam optimizer with an initial learning rate of 0.00005, halved every 3 epochs. Each training batch contains 120 places, with 4 images per place (i.e., 480 images). Besides, we set the maximum epochs to 20.
@@ 4.3 Comparisons with State-of-the-Art Methods
This section shows the experimental comparison of our ImAge with SOTA methods, including 11 single-stage VPR methods: NetVLAD Arandjelovic et al. (2016), SFRS Ge et al. (2020), CosPlace Berton et al. (2022a), MixVPR Ali-Bey et al. (2023), EigenPlaces Berton et al. (2023), CricaVPR Lu et al. (2024b), SALAD Izquierdo and Civera (2024b), SALAD-CM Izquierdo and Civera (2024a), BoQ Ali-bey et al. (2024), SuperVLAD Lu et al. (2024d) and EDTformer Jin et al. (2025), as well as 2 two-stage VPR methods (TransVPR Wang et al. (2022) and SelaVPR Lu et al. (2024c)) that leverage local features for re-ranking. The latest studies, CricaVPR, SALAD, SALAD-CM, BoQ, SelaVPR, SuperVLAD, and EDTformer, all use the foundation model DINOv2 as the backbone to extract deep features and achieve SOTA results. Our method mainly adopts DINOv2-base-register in experiments. Additionally, Cosplace and EigenPlaces construct an extra large-scale dataset (SF-XL) for training. CircaVPR, SALAD, BoQ, and EDTformer are trained on the GSV-Cities dataset, while SALAD-CM combines GSV-Cities and MSLS-train for training. Our work further merges Pitts30k-train, MSLS-train, SF-XL, and GSV-Cities for training, following the process in SelaVPR++ Lu et al. (2025). Table 2 presents the comprehensive quantitative results. Moreover, to enable a fairer comparison among three leading aggregation methods (NetVLAD, SALAD, and BoQ) and our ImAge, we conduct a consistent comparison using the same setup (backbone, training data, image resolution), as shown in Table 3. The experiments using other transformer backbones (ViT and CLIP) are shown in Appendix D.
@@
NetVLAD Arandjelovic et al. (2016) 32768 81.9 91.2 93.7 53.1 66.5 71.1 35.1 47.4 51.7 60.6 68.9 74.6 6.4 10.1 12.5
SFRS Ge et al. (2020) 4096 89.4 94.7 95.9 69.2 80.3 83.1 41.6 52.0 56.3 81.0 88.3 92.4 16.1 23.9 28.4
TransVPR Wang et al. (2022) / 89.0 94.9 96.2 86.8 91.2 92.4 63.9 74.0 77.5 79.0 82.2 85.1 63.5 68.5 70.2
CosPlace Berton et al. (2022a) 512 88.4 94.5 95.7 82.8 89.7 92.0 61.4 72.0 76.6 81.9 90.2 92.7 58.5 73.7 79.4
MixVPR Ali-Bey et al. (2023) 4096 91.5 95.5 96.3 88.0 92.7 94.6 64.0 75.9 80.6 85.1 91.7 94.3 76.2 86.9 90.3
EigenPlaces Berton et al. (2023) 2048 92.5 96.8 97.6 89.1 93.8 95.0 67.4 77.1 81.7 93.0 96.2 97.5 71.2 83.8 88.1
SelaVPR Lu et al. (2024c) / 92.8 96.8 97.7 90.8 96.4 97.2 73.5 87.5 90.6 94.0 96.8 97.5 87.3 93.8 95.6
CricaVPR Lu et al. (2024b) 4096 94.9 97.3 98.2 90.0 95.4 96.4 69.0 82.1 85.7 93.0 97.5 98.1 90.7 96.3 97.6
SuperVLAD Lu et al. (2024d) 3072 95.0 97.4 98.2 92.2 96.6 97.4 75.3 86.8 89.9 95.2 97.8 98.1 91.0 96.4 97.7
SALAD Izquierdo and Civera (2024b) 8448 92.5 96.4 97.5 92.2 96.4 97.0 75.0 88.8 91.3 94.6 97.5 97.8 89.7 95.5 97.0
SALAD-CM Izquierdo and Civera (2024a) 8448 92.7 96.8 97.9 94.2 97.2 97.4 82.7 91.2 92.7 96.8 97.5 97.8 96.0 98.5 99.2
BoQ Ali-bey et al. (2024) 12288 93.7 97.1 97.9 93.8 96.8 97.0 79.0 90.3 92.0 96.5 97.8 98.4 90.6 96.0 97.5
EDTformer Jin et al. (2025) 4096 93.4 97.0 97.9 92.0 96.6 97.2 78.4 89.8 91.9 97.1 98.1 98.4 88.3 95.3 97.0
@@
Besides, we also conduct the consistent comparison experiments on some supplementary datasets, including Baidu Mall Sun et al. (2017), SPED Chen et al. (2017a), Pitts250k Torii et al. (2013), St. Lucia Glover et al. (2010), and SVOX Berton et al. (2021b), and the results are shown in Table 4. Compared to the three SOTA explicit aggregation methods, our ImAge achieves the best R@1 performance on all supplementary datasets. In particular, on Baidu Mall, which is the only indoor dataset and exhibits a distinct visual distribution from the other outdoor datasets, our method achieves the best performance, outperforming NetVLAD, SALAD, and BoQ with 0.8%, 3.3%, and 5.0% absolute R@1 improvements, respectively. This demonstrates that the global descriptors produced by our ImAge method through implicit aggregation are not only highly robust against common visual changes but also exhibit superior generalization ability.
@@ 4.4 Ablation Studies
Effect of aggregation tokens initialization. To validate the effectiveness of our proposed initialization methods for agg tokens, we conduct an ablation study using four initialization strategies: zero initialization (i.e., no initialization), normal distribution initialization (commonly used for the class token or register tokens initialization Darcet et al. (2024)), vanilla cluster centers (yielded by $k$-means) initialization, and L2-normalized cluster centers initialization (i.e., ours). We consistently use 8 agg tokens and prepend them to the patch tokens before the fourth-to-last transformer block. The experimental results are presented in Table 6. Zero initialization produces uniform representations at the beginning, limiting (even harming) interaction between agg tokens and patch features, and hindering global context modeling. In contrast, normal initialization provides a better inductive bias during early training and introduces slight randomness into the agg tokens, which helps break symmetry to get better performance. However, both initialization methods lack any visual prior, forcing the agg tokens to learn the patterns relevant to VPR from scratch, which constrains their final performance. Initializing agg tokens with cluster centers can be viewed as injecting a data-driven prior. These centers, obtained via unsupervised clustering of descriptors from randomly sampled training images, tend to capture common visual patterns. Such initialization can facilitate agg tokens to learn meaningful global information and diminish useless elements. Moreover, L2-normalized cluster centers offer more robust initializations for agg tokens by mitigating the influence of outliers, thereby achieving the optimal performance on all datasets.
@@ Appendix B Limitations and Future Work
While our study provides a novel insight (i.e., implicit aggregation without an explicit aggregator) into achieving robust global image representation for VPR, we acknowledge three limitations of our work: 1) Although our ImAge method demonstrates universality for the transformer-based models, compared to the foundation models pre-trained on the massive dataset (e.g., DINOv2 pre-trained on the LVD-142M dataset Oquab et al. (2023)), the performance improvement is less pronounced when using backbones without pre-training on sufficiently massive data (e.g., the ViT pre-trained only on ImageNet Deng et al. (2009)). This will be shown in Appendix D. However, this also suggests that with the advancement of increasingly powerful foundation models, the superiority of our approach compared to existing VPR methods may become more prominent. 2) The proposed method may not be a good choice when we want to keep the backbone frozen (like AnyLoc Keetha et al. (2023)) or when the backbone is extremely expensive to fine-tune. However, it is also worth noting that we only fine-tune the last few blocks of the transformer backbone in most cases, which is relatively cheap. 3) Although this work focuses on the VPR task, we believe that the proposed ImAge is broadly applicable to a wide range of image (or other modalities) retrieval tasks. The potential of our ImAge method for more applications in the machine learning community needs to be further explored through more experiments in future work.
@@ Appendix C More Details about the Relations & Differences to Other Methods
Our ImAge design draws inspiration from prior works, particularly DINOv2-register Darcet et al. (2024) and prompt tuning Lester et al. (2021). Both approaches introduce additional tokens to the transformer backbone before the first encoder block. However, their objectives and usage differ fundamentally from ours. Prompt Tuning aims to adapt a frozen model to downstream tasks by learning a set of prompt tokens in a parameter-efficient manner, while these tokens are typically excluded from the final representation. DINOv2-register introduces additional register tokens to mitigate artifacts in the feature maps. Although the study shows that these registers may capture certain global information, they are ultimately discarded, and only the patch and class tokens are used for downstream tasks. In contrast, our ImAge method introduces the aggregation (agg) tokens before a particular transformer block, and utilizes agg tokens from the output of the last block as the final global image representation, which provides a reverse perspective compared to these approaches. In addition, while the class token within the transformer backbone is sometimes used as a global representation, our ImAge differs in several key aspects and demonstrates superior scalability. First, the class token is typically introduced at the beginning of the transformer and starts to learn from shallow features, which may limit its flexibility and make it difficult to fully capture task-specific complex semantics. Second, the class token is a single fixed embedding, which inherently restricts its representational capacity. Although it may suffice for relatively simple classification tasks, it often proves inadequate for more complex scenarios requiring richer and more flexible representations. In contrast, our ImAge introduces agg tokens with customizable positions and quantities, allowing them to fully learn task-specific global features. Moreover, we also design a tokens initialization method based on the $k$-means algorithm, which is significantly different from other works. Additionally, among existing VPR methods, BoQ Ali-bey et al. (2024) also offers some valuable insights for our work. However, it elaborately designs an explicit aggregator consisting of additional encoder blocks and cross-attention layers, which aggregates global information from patch tokens into a set of extra learnable queries. In contrast, we focus on the transformer backbone itself and make use of the inherent self-attention mechanism. Our study reveals a new insight: the aggregation function, previously implemented through an exquisitely designed aggregator, already appears naturally in the transformer backbone. We demonstrate that, by adding just some additional tokens, we can fully develop this implicit and progressive aggregation behavior.
@@ Appendix D Comparison to NetVLAD Using Other Transformer Backbones
In the main paper, we conduct experiments using the DINOv2 backbone to validate the effectiveness of our method. Notably, DINOv2 is a foundation model based on the ViT architecture and pre-trained on the large-scale curated LVD-142M dataset. However, our method is also applicable to other transformer models. To this end, we conduct additional experiments using the CLIP Radford et al. (2021) model and a ViT model pre-trained only on ImageNet. The results are shown in Table 8. We observe that our proposed method, ImAge, consistently achieves higher R@1 performance across all datasets compared to the explicit aggregation method NetVLAD. However, the performance gains of ImAge are less pronounced (except for ViT-ImAge on MSLS-val) compared to using DINOv2-base-register as the backbone. This observation aligns with the prior study Kerssies et al. (2025), which suggests that foundation models pre-trained on large-scale datasets (significantly larger than ImageNet) are more capable of utilizing additional tokens to capture global information. Moreover, using CLIP as the backbone yields significantly less improvement than DINOv2. Although CLIP is a widely used foundation model (i.e., a vision-language model), its pre-training data and objectives differ considerably from those of the VPR task, making it not a promising choice. This is consistent with the prior work AnyLoc Keetha et al. (2023), which suggests that CLIP performs significantly worse than DINOv2 in outdoor VPR scenarios.
Table 8: Results of NetVLAD and our ImAge using CLIP (base version, only vision encoder) and ViT (base version) as backbone. All models are trained on the GSV-Cities dataset with the batch size equal to 120. The learning rate is 0.00006 for the CLIP-based model and 0.0003 for the ViT-based model. For ViT, the last two blocks are directly truncated and all other blocks are trainable, as in Berton et al. (2022b); Lu et al. (2024d). For CLIP, we only train the last 6 blocks with the previous layers frozen. All methods produce 768*8-dimensional descriptors, i.e., 8 clusters for NetVLAD and 8 aggregation tokens for ImAge, the same as in the main paper.
@@ Appendix E Improving Other VPR Methods with ImAge
Since our ImAge is a general image representation method for VPR, it can not only be implemented based on different transformer backbones, but also can be combined with some other VPR methods to improve their performance. This section uses the CricaVPR Lu et al. (2024b) method as an example to conduct experiments, and the results are shown in Table 9. It can be seen that our method significantly improves the performance.
@@ Appendix I More Details about Datasets
The testing datasets used in our experiments, including Pitts30k, Pitts250k, Tokyo24/7, Nordland, SPED, St. Lucia, and SVOX, are organized following the Visual Geo-localization (VG) benchmark Berton et al. (2022b). Notably, we use the official version MSLS dataset as in previous work Warburg et al. (2020); Ali-Bey et al. (2023); Lu et al. (2024b); Ali-bey et al. (2024); Izquierdo and Civera (2024b). This version of MSLS-val only consists of 740 query images, which is different from the version in the VG benchmark Berton et al. (2022b). In addition, there are also several versions of the Nordland dataset in the VPR task. In our experiments, we use the version in the VG benchmark Berton et al. (2022b), which employs the summer images as the database and the winter images as queries, each containing 27592 images. Baidu Mall Sun et al. (2017) is a well-known indoor dataset for image-based localization. All images are collected at a shopping mall that is over 5000 square meters with many challenging elements, such as transparent windows, reflective materials, repetitive structures, dynamic pedestrians, etc.
Moreover, in the comprehensive comparison (i.e., Table 2 in main paper) with other SOTA methods, we merge Pitts30k-train, MSLS-train, SF-XL, and GSV-Cities for training, following the approach in SelaVPR++ Lu et al. (2025). Specifically, we process datasets other than GSV-Cities to divide places into a finite number of categories, thus facilitating fully supervised training with the multi-similarity loss Wang et al. (2019).
@@
NetVLAD Arandjelovic et al. (2016) and SFRS Ge et al. (2020) both consist of a VGG16 backbone and a NetVLAD aggregator, and use Pitts30k as the training dataset. The latter employs self-supervised image-to-region similarities to mine hard positive samples for training a more robust model. In our experiments, we use their PyTorch implementations for comparison.
CosPlace Berton et al. (2022a) and EigenPlaces Berton et al. (2023) both frame VPR training as a classification task and use the SF-XL dataset to train their models. For Cosplace, we use the official model based on the VGG16 backbone (with the 512-dim output feature) for testing. For EigenPlaces, we utilize its official implementation and the best configuration based on the ResNet50 backbone to output 2048-dim global descriptors for comparison.
MixVPR Ali-Bey et al. (2023) aggregates the deep features using the multi-layer perceptrons and trains the model with multi-similarity loss Wang et al. (2019) on the GSV-Cities Ali-bey et al. (2022) dataset. We apply its best-performing configuration (ResNet50 with 4096-dim output features) for comparison.
CricaVPR Lu et al. (2024b), SuperVLAD Lu et al. (2024d), SALAD Izquierdo and Civera (2024b), BoQ Ali-bey et al. (2024), and EDTformer Jin et al. (2025) all use the foundation model DINOv2-base Oquab et al. (2023) as the backbone to extract deep features, and train their models on GSV-Cities with the multi-similarity loss. In the comparison experiments, we consistently use their official implementations and the best configurations.
SALAD-CM Izquierdo and Civera (2024a) is an improvement of SALAD. This work analyzes the Geographic Distance Sensitivity of VPR embeddings and proposes a novel mining strategy to address it. Moreover, SALAD-CM first trains the model using both the GSV-Cities and MSLS datasets for better performance. In the comparison experiments, we follow its official implementation.
The rest TransVPR Wang et al. (2022) and SelaVPR Lu et al. (2024c) are two-stage VPR methods. These works provide two models: one trained for testing on urban datasets (e.g., Pitts30k and Tokyo24/7), and another trained for testing on datasets that may contain suburban and natural scenes (e.g., MSLS and Nordland). We follow the usage in their original paper for comparison experiments.
@@
Fig. 8 is the snapshot of the MSLS place recognition challenge Warburg et al. (2020) leaderboard at the time of submission, and our ImAge method ranks 1st.